# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ 8 Conclusion
In this work, we aimed at establishing a connection between notions of uncertainty coming from information theory and conformal prediction. We proved that we can use split conformal prediction methods to upper bound the conditional entropy of the target variable given the inputs, and that these upper bounds also form principled objectives for conformal training. We empirically validated our approach to conformal training, with strong results in both centralized and federated settings. Unfortunately, our experiments have not singled out the best performing among our bounds. The general trend we observe is that the DPI and MB Fano bounds perform better on complex classification tasks, while simple Fano seems to excel in relatively simple tasks or where there is a higher risk of overfitting, like in the federated setting.
@@
Further, the information theoretic perspective also offers a simple yet rigorous approach to incorporate side information into conformal prediction, which we experimentally demonstrate leads to better predictive efficiency in almost all cases. To the best of our knowledge this is the first attempt at connecting information theory and conformal prediction. Given the thus far limited communication between these two research communities, we expect our work to incite a fruitful exchange of not only ideas but also theory and algorithms between these two research domains. We see an extension of our methods to the regression setting, as a particularly promising avenue for future work.
Appendix A Background on Conformal Prediction
@@
In this section we outline a brief introduction to conformal prediction, providing the reader with the necessary background to follow our main results. Readers already familiar with conformal prediction can safely skip this section. We start by reviewing the definitions of quantiles and exchangeability, which are central to the main results in conformal prediction.
A.1 Exchangeability, Ranks and Quantiles
@@
The main assumption in conformal prediction is that the data points used for calibration and testing are exchangeable. Next, we define the concept of exchangeability and discuss how it leads to the main results in conformal prediction via properties of ranks of exchangeable random variables. Our exposition is markedly brief, and we refer the reader to Kuchibhotla (2020) for a more thorough discussion on exchangeability and its importance in conformal prediction.
Formally, the concept of exchangeability can be defined as below.
Definition A.1 (Exchangeable Random Variables).
@@
Note that exchangeability is a weaker assumption than the i.i.d. (independent and identically distributed) assumption commonly relied upon in machine learning. More precisely, exchangeable random variables must be identically distributed but not necessarily independent Kuchibhotla (2020). Naturally, i.i.d. random variables are also exchangeable.
One relevant consequence of exchangeability, that is central to conformal prediction, is that the ranks of exchangeable random variables are uniformly distributed. We define ranks and this property more formally in Definition A.2 and Lemma A.4, respectively.
Definition A.2 (Rank).
@@
In words, Lemma A.4 tell us that the ranking of exchangeable random variables is uniformly distributed among all possible permutations $\pi:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}.$ That means the probability of observing any one ranking is equal to $\nicefrac{{1}}{{n!}}$ and, importantly, independent of the distribution of $X$. The corollary below follows directly from Lemma A.4.
Corollary A.1.
@@
As we shall see, Corollary A.1 is central to the main result in conformal prediction, but before proving that result, we should first define the concept of quantiles.
Definition A.5 (Quantile).
@@ A.2 Conformal Prediction
where the probability is over the randomness of $\{(X_{i},Y_{i})\}_{i=1}^{n}\cup\{(X_{test},Y_{test})\}$. In words, this guarantees us that the correct label $Y_{test}$ is included in the constructed set ${\mathcal{C}}(X_{test})$ with probability at least $1-\alpha.$
There are different ways to achieve this guarantee, but for simplicity we will focus on split conformal prediction or SCP Papadopoulos et al. (2002), since it is easier to grasp than other variants of conformal prediction and is also the main object of study in this paper. We start by assuming a calibration dataset ${\mathcal{D}}_{cal}$ which consists of $n$ i.i.d. samples $(X_{i},Y_{i})$ drawn from an unknown distribution over ${\mathcal{X}}\times{\mathcal{Y}}$. We also assume access to a model $f:{\mathcal{X}}\to\hat{{\mathcal{Y}}}$, where the output space $\hat{{\mathcal{Y}}}$ can be different from ${\mathcal{Y}}$. Prediction sets satisfying the guarantee above can be constructed via the following three steps.
@@
(same result as in the main text (Vovk et al. (2005); Lei et al. (2018))
@@
For simplicity we assume that the set of scores $\{S_{i}\}_{i=1}^{n}$ are distinct (or have been made distinct by a suitable random tie-breaking rule). Denote $S_{test}=s(X_{test},Y_{test})$ and observe that since $s(\cdot)$ is applied element-wise to each pair $(X_{i},Y_{i})$ it preserves exchangeability, and thus random variables $\{S_{i}\}_{i=1}^{n}\cup\{S_{test}\}$ are also exchangeable. Next, we show that the following events are all equivalent
@@
Finally, this where the crucial exchangeability assumption comes into play, allowing us to apply Corollary A.1 to get
@@
From there it is easy to verify that the right hand side is at least $1-\alpha$ and at most $1-\alpha+\nicefrac{{1}}{{n+1}}.$
Appendix B Background on List Decoding and Basic Information Theoretic Results
In this section we enunciate some classical results from information theory that are instrumental in deriving our main theoretical results, as we show in the detailed proofs in Appendix C. We also provide a brief introduction to list decoding and demonstrate conformal prediction can be framed as a list decoding problem.
B.1 Data processing inequalities for f-divergence
@@
Definition B.2 (Conditional $f$-Divergence).
@@
Theorem B.3 is the classical data processing inequality (DPI), which we restate below for the sake of completeness. The classical versions of DPI, stated in terms of mutual information, can be found in classical information theoretic text books like Cover & Thomas (2006), while the generalization of DPI to $f$-divergence can be found in other works with a comprehensive survey in Polyanskiy & Wu (2014); Sason & Verdú (2016).
Theorem B.3 (Data Processing Inequality for $f$-divergence).
@@
The proof can be found in standard textbooks in information theory, see for example Chapter 7, Section 7.2. in Polyanskiy & Wu (2014) for a derivation of the DPI enunciated in the same form as above.
@@
Theorem B.4 (Data Processing Inequality for Conformal Prediction).
@@
Theorem B.5 (Conditional Data Processing Inequality for Conformal Prediction).
@@
and from the monotonicity property of the f-divergence (Polyanskiy & Wu, 2014) we have that
@@
B.2 List Decoding
List decoding Elias (1957) is a notion coming from coding theory, a large branch of engineering and mathematics that arises from the application of information theory to the design of reliable communication systems and robust information processing and storage. In particular, we are interested in channel coding, a field concentrated on the design of so-called error-correcting codes to enable reliable communication over inaccurate or noisy communication channels. The general setup studied in channel coding, including list decoding, can be summarized as follows. A message $y\in{\mathcal{Y}}$ is encoded and transmitted over a noisy channel, and a message $\hat{y}\in\hat{Y}$ is received. The noisy channel is governed by probability density $p(\hat{y}|y)$ that describes the probability of observing output $\hat{Y}\in\hat{\mathcal{Y}}$ given input $Y\in{\mathcal{Y}}$. The receiver attempts to decode $\hat{y}$, that is, to guess the originally transmitted message $y$ from the received one, $\hat{y}$. At this point, parallels to machine learning should already have become clear. If the receiver provides a single guess of the transmitted message $y$, we are in a unique-decoding scenario, which is akin to a point prediction in machine learning. Conversely, if the receiver is allowed to guess a list of the most likely messages, we have list-decoding, which closely resembles conformal prediction.
More formally, a list-decoding algorithm can be defined by a set predictor $L:\hat{\mathcal{Y}}\rightarrow 2^{{\mathcal{Y}}}$, with maximum output set size $|L(\hat{Y})|\leq M.$ Naturally, the goal is to design the function $L$ so as to maximize the probability of $Y\in L(\hat{Y}).$ Similarly, for a given input-label pair $(X,Y)$ the goal of conformal prediction is to provide a set based on the output of a given model $\hat{Y}=f(X)$ that contains $Y$ with a pre-determined confidence. For instance, $\hat{Y}$ could be the logits output by the model, i.e., a vector with the dimension equal to number of classes. If we consider $\hat{Y}$ the noisy observation of the ground-truth $Y$, then the problem of determining a set containing $Y$ is the list decoding problem. Since $\hat{Y}=f(X)$ is a function of the input data $X$, we can consider the communication channel $\mathbb{P}(X|Y)$, which motivates our bounds on the conditional entropy $H(Y|X).$ Alternatively, we can also consider the communication channel $\mathbb{P}(\hat{Y}|Y;X)$ directly, which justifies applying the same upper bounds to $H(Y|\hat{Y}).$ Note that in this last case, the input data $X$ can also be taken as side information, although it is rarely used directly in building the conformal prediction set.
@@ B.2 List Decoding
B.3 List decoding: Information Theoretic Inequalities
Fano’s inequality for variable size list decoding. The following generalization of Fano’s inequality is given in [Raginsky & Sason (2013), Appendix 3.E].
Theorem B.6.
@@ B.3 List decoding: Information Theoretic Inequalities
It is easy to see that the above set is the smallest set given each $y$ and the confidence level $1-\alpha$ (see, for example, Merhav (2014)). The same result holds in conformal prediction Vovk et al. (2016); Romano et al. (2020).
B.4 Fano and Data Processing Inequalities for Conformal Prediction
First, from Fano’s inequality for list decoding, Theorem B.6, we get the next result ‘out-of-the-box’.
Proposition B.7.
@@ B.4 Fano and Data Processing Inequalities for Conformal Prediction
The proposition follows easily from Theorem B.6 by using the condition $\mathbb{P}(Y\in{\mathcal{C}}(x))\geq 1-\alpha$. Note that in Theorem B.6, we assume non-empty lists, whereas in Proposition B.7 we allow empty prediction sets but apply the maximum operator $[x]^{+}:=\max\{x,0\}$. This justified because the last term of Fano’s inequality relates to the probability of correct assignments $Y\in{\mathcal{C}}(X)$ which never happens for empty sets. See Proposition D.1 for the proof, where the same result appears.
@@
The bounds that we present in the main paper, model-based and simple Fano bounds, are actually derived through a slightly different root by leveraging the lower and upper bounds in the finite-sample guarantee of conformal prediction (Theorem 2.1). We derive these other two bounds in Appendix C.2.
B.5 Related Work on Information Theoretic Inequalities
@@
The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s—see for instance (Ali & Silvey, 1966; Csiszár, 1967; Morimoto, 1963). A key $f$-divergence inequality is the data processing inequality, which was used in information theory to establish various upper bounds on the achievability of coding schemes for different tasks—see (Sason & Verdú, 2016; Sason, 2019) for an extensive survey. Moreover, the data processing inequality for $f$-divergences provides a unified way of obtaining many classical and new results, for example Fano’s inequality (Fano, 1949). The tightness of data processing inequalities is discussed in terms of Bregman’s divergence in (Collet, 2019; Liese & Vajda, 2006) and in terms of $\chi^{2}$-divergence in Sason (2019).
@@
When it comes to list decoding, there are a number of relevant inequalities in the literature. List decoding was introduced in context communication design by Elias (Elias, 1957) and Wozencraft (Wozencraft, 1958). For fixed list size, the information theoretic bounds on list decoding were obtained in (Gallager, 1986) using error exponent analysis. A generalization of Fano’s inequality to list decoding was given in (Csiszár & Körner, 2011) in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in (Raginsky & Sason, 2013). For fixed list size, stronger inequalities, some based Arimoto-Rényi conditional entropy, were presented in (Sason & Verdú, 2017). Variable size list decoding was discussed in (Sason, 2019) using the notion of $E_{\gamma}$ resolvability first introduced in (Polyanskiy et al., 2010) related to the dependence testing bound. It was used again in context of channel resolvability in (Liu et al., 2017), where some relevant inequalities have been obtained and discussed. A selection of most relevant inequalities for list decoding can be found in (Sason, 2019).
Appendix C Proofs of Main Theoretical Results

### v2

@@ 8 Conclusion
In this work, we established a link between notions of uncertainty coming from conformal prediction and information theory (or more precisely variable-size list decoding). We proved that one can use split conformal prediction methods to upper bound the conditional entropy of the target variable given the inputs, and that these upper bounds form principled objectives for conformal training. We empirically validated our approach to conformal training, with strong results in both centralized and federated settings. Furthermore, the information-theoretic perspective also offers a simple yet rigorous approach to incorporate side information into conformal prediction, which we experimentally show leads to better predictive efficiency. To the best of our knowledge, this is the first attempt at connecting information theory and conformal prediction. Given the limited communication between these two research communities thus far, we expect our work to incite a fruitful exchange of not only ideas but also theory and algorithms between these two research domains. We see an extension of our methods to the regression setting, as a particularly promising avenue for future work.
Appendix A Broader Impact, Limitations and Other Remarks
Broader Impact.
This work explores the connection between conformal prediction and information theory, with the end goal of advancing the current state of the art of uncertainty estimates in machine learning. With that in mind, we believe its potential societal consequences are chiefly positive, since our work might contribute to a larger adoption of uncertainty estimates, in a number of safety-critical applications. Notwithstanding, conformal prediction, like any other uncertainty estimation method, should be applied with care and a proper understanding of the provided guarantees so as not to create an illusion of safety when there is none. That is why our work aims to develop new methods and algorithms from first principles, with the hope of providing new techniques that we can study and understand at a deeper level so as to mitigate, or at least foresee, some of these failure modes.
Limitations.
Our work is not without limitations. In practice, estimating the conditional entropy $H(Y|X)$ is a difficult problem. That is part of the reason upper bounds might be useful, but that also means it is hard to evaluate how tight our upper bounds are. This also limits potential applications of our upper bounds like, for instance, estimating the expected prediction set size. Our experiments in Appendix G.3 rely on quantization to get a reasonable lower bound on the conditional entropy, and thus lower bound the expected prediction set size. Finally, in terms of our experimental results, we unfortunately have not been able to single out which of our new upper bounds performs best. The general trend we observe is that the DPI and model-based Fano bounds perform better on complex classification tasks, while simple Fano seems to excel in relatively simple tasks or where there is a higher risk of overfitting, like in the federated setting.
Regression Setting.
In this paper, we focus our experiments on the classification setting, similarly to previous works on conformal training [8, 59]. However, in principle, our bounds pose no assumptions on the underlying prediction problem and we see the application of our results to the regression setting as a promise avenue for future work. In practice, only the simple Fano bound in (6) would not be directly applicable to the regression setting, but that is mainly because it assumes a uniform distribution over the output space and we get the, potentially infinite, $|\mathcal{Y}|$ term. The other two bounds we propose, DPI in (4) and model-based Fano in (5), do not include the $|\mathcal{Y}|$ and $|C(X)|$ terms and can be applied to regression problems as is.
Computational Cost.
The computational cost of using our bounds for conformal training is the same as that of previously proposed conformal training [59]. In comparison to regular training, i.e. minimizing the cross-entropy loss, the additional cost is given by the (differentiable) sorting operation of the scores, which in our case was performed with diffsort [48] with bitonic networks, which has complexity $\mathcal{O}(b^{2}\log b^{2})$ for $b$ the batch size. Since the batch size is typically small, the additional computation cost is only marginal in practice.
Appendix B Background on Conformal Prediction
@@
In this section, we outline a brief introduction to conformal prediction, providing the reader with the necessary background to follow our main results. Readers already familiar with conformal prediction can safely skip this section. We start by reviewing the definitions of quantiles and exchangeability, which are central to the main results in conformal prediction.
B.1 Exchangeability, Ranks and Quantiles
@@
The main assumption in conformal prediction is that the data points used for calibration and testing are exchangeable. Next, we define the concept of exchangeability and discuss how it leads to the main results in conformal prediction via properties of ranks of exchangeable random variables. Our exposition is markedly brief, and we refer the reader to [24] for a more thorough discussion on exchangeability and its importance in conformal prediction.
Formally, the concept of exchangeability can be defined as follows:
Definition B.1 (Exchangeable Random Variables).
@@
Note that exchangeability is a weaker assumption than the i.i.d. (independent and identically distributed) assumption commonly relied upon in machine learning. More precisely, exchangeable random variables must be identically distributed but not necessarily independent [24]. Naturally, i.i.d. random variables are also exchangeable.
One relevant consequence of exchangeability that is central to conformal prediction is that the ranks of exchangeable random variables are uniformly distributed. We define ranks and this property more formally in Definition B.2 and Lemma B.4, respectively.
Definition B.2 (Rank).
@@
In words, Lemma B.4 tell us that the ranking of exchangeable random variables is uniformly distributed among all possible permutations $\pi:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}.$ That means the probability of observing any one ranking is equal to $\nicefrac{{1}}{{n!}}$ and, importantly, independent of the distribution of $X$. The corollary below follows directly from Lemma B.4.
Corollary B.1.
@@
As we shall see, Corollary B.1 is central to the main result in conformal prediction, but before proving that result, we should first define the concept of quantiles.
Definition B.5 (Quantile).
@@ B.2 Conformal Prediction
where the probability is over the randomness of $\{(X_{i},Y_{i})\}_{i=1}^{n}\cup\{(X_{test},Y_{test})\}$. Since the probability is taken over both $X_{test}$ and $Y_{test}$, this means that on average across all possible values in ${\mathcal{X}}$, the correct label $Y_{test}$ is included in the constructed set ${\mathcal{C}}(X_{test})$ with probability at least $1-\alpha.$ This property is known as marginal coverage. This is in contrast to the stronger conditional coverage, which requires
$\mathbb{P}(Y_{test}\in{\mathcal{C}}(X_{test})|X_{test})\geq 1-\alpha.$
That is, the guarantee holds for each $X_{test}$ individually, with $X_{test}$ fixed and the probability taken over the randomness of $Y_{test}$ only. However, distribution-free conditional coverage is known to admit no non-trivial solution (i.e., besides ${\mathcal{C}}(X_{test})={\mathcal{Y}}$) [61, 28]. Therefore, in this paper, whenever we refer to the lower and upper bounds provided by conformal prediction, we mean those provided by Theorem 2.1, which guarantees only marginal coverage.
There are different ways to achieve marginal coverage, but for simplicity we will focus on split conformal prediction or SCP [45], since it is easier to grasp than other variants of conformal prediction and is also the main object of study in this paper. We start by assuming a calibration dataset ${\mathcal{D}}_{cal}$ which consists of $n$ i.i.d. samples $(X_{i},Y_{i})$ drawn from an unknown distribution over ${\mathcal{X}}\times{\mathcal{Y}}$. We also assume access to a model $f:{\mathcal{X}}\to\hat{{\mathcal{Y}}}$, where the output space $\hat{{\mathcal{Y}}}$ can be different from ${\mathcal{Y}}$. Prediction sets satisfying the guarantee above can be constructed via the following three steps.
@@
(same result as in the main text (Vovk et al. [62], Lei et al. [29])
@@
For simplicity, we assume that the set of scores $\{S_{i}\}_{i=1}^{n}$ are distinct (or have been made distinct by a suitable random tie-breaking rule). Denote $S_{test}=s(X_{test},Y_{test})$ and observe that since $s(\cdot)$ is applied element-wise to each pair $(X_{i},Y_{i})$ it preserves exchangeability, and thus random variables $\{S_{i}\}_{i=1}^{n}\cup\{S_{test}\}$ are also exchangeable. Next, we show that the following events are all equivalent
@@
Finally, this is where the crucial exchangeability assumption comes into play, allowing us to apply Corollary B.1 to get
@@
From there, it is easy to verify that the right hand side is at least $1-\alpha$ and at most $1-\alpha+\nicefrac{{1}}{{n+1}}.$
Appendix C Background on List Decoding and Basic Information Theoretic Results
In this section, we enunciate some classical results from information theory that are instrumental in deriving our main theoretical results, as we show in the detailed proofs in Appendix D. We also provide a brief introduction to list decoding and demonstrate conformal prediction can be framed as a list decoding problem.
C.1 Data processing inequalities for f-divergence
@@
Definition C.2 (Conditional $f$-Divergence).
@@
Theorem C.3 is the classical data processing inequality (DPI), which we restate below for the sake of completeness. The classical versions of DPI, stated in terms of mutual information, can be found in classical information theoretic text books like [13], while the generalization of DPI to $f$-divergence can be found in other works with a comprehensive survey in [49, 56].
Theorem C.3 (Data Processing Inequality for $f$-divergence).
@@
The proof can be found in standard textbooks in information theory see, for example, Chapter 7, Section 7.2. in [49] for a derivation of the DPI enunciated in the same form as above.
@@
Theorem C.4 (Data Processing Inequality for Conformal Prediction).
@@
Theorem C.5 (Conditional Data Processing Inequality for Conformal Prediction).
@@
and from the monotonicity property of the f-divergence [49] we have that
@@
C.2 List Decoding
List decoding [17, 66] is a notion coming from coding theory, a large branch of engineering and mathematics that arises from the application of information theory to the design of reliable communication systems and robust information processing and storage. In particular, we are interested in channel coding, a field concentrated on the design of so-called error-correcting codes to enable reliable communication over inaccurate or noisy communication channels.
The general setup studied in channel coding, including list decoding, can be summarized as follows. A message $y\in{\mathcal{Y}}$ is encoded and transmitted over a noisy channel, and a message $x\in{\mathcal{X}}$ is received. The noisy channel is governed by probability density $p(x|y)$ that describes the probability of observing output $x\in{\mathcal{X}}$ given input $y\in{\mathcal{Y}}$. The receiver attempts to decode $x$, that is, to guess the originally transmitted message $y$ from the received one, $x$. At this point, parallels to machine learning should already have become clear. If the receiver provides a single guess of the transmitted message $y$, we are in a unique-decoding scenario, which is akin to a point prediction in machine learning. Conversely, if the receiver is allowed to guess a set (or a list in the list decoding formalism) of the most likely messages, we have list decoding, which closely resembles conformal prediction. Note that in many settings, a list decoding algorithm is constrained to output a list of fixed size. While simple conformal prediction methods for regression settings show the same limitation, the parallel between the two domains is more pertinent when we consider variable-size list decoding [55].
More formally, a list-decoding algorithm can be defined by a set predictor $L:{\mathcal{X}}\rightarrow 2^{{\mathcal{Y}}}$, with maximum output set size $|L(X)|\leq M.$ Naturally, the goal is to design the function $L$ so as to maximize the probability of $Y\in L(X).$ Similarly, for a given input-label pair $(X,Y)$ the goal of conformal prediction is to provide a set that contains $Y$ with a certain pre-determined probability. It is this connection that motivates our bounds on the conditional entropy $H(Y|X).$ However, the nonconformity score in conformal prediction is typically a function of the output of a given model $f:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. For instance, $\hat{Y}=f(X)$ could be the logits output by the model, i.e., a vector with the dimension equal to number of classes. If we consider $\hat{Y}$ the noisy observation of the ground-truth $Y$, then the problem of determining a set containing $Y$ is again the list decoding problem. Therefore, we can also consider the communication channel $p(\hat{y}|y;x)$ directly, which justifies applying the same upper bounds to $H(Y|\hat{Y}).$ Note that in this last case, the input data $X$ can also be taken as side information, although it is rarely used directly in building the conformal prediction set.
@@ C.2 List Decoding
C.3 List decoding: Information Theoretic Inequalities
Fano’s inequality for variable size list decoding. The following generalization of Fano’s inequality is given in [[51], Appendix 3.E].
Theorem C.6.
@@ C.3 List decoding: Information Theoretic Inequalities
It is easy to see that the above set is the smallest set given each $y$ and the confidence level $1-\alpha$ (see, for example, [41]). The same result holds in conformal prediction [63, 53].
C.4 Fano and Data Processing Inequalities for Conformal Prediction
First, from Fano’s inequality for list decoding, Theorem C.6, we get the next result ‘out-of-the-box’.
Proposition C.7.
@@ C.4 Fano and Data Processing Inequalities for Conformal Prediction
The proposition follows easily from Theorem C.6 by using the condition $\mathbb{P}(Y\in{\mathcal{C}}(x))\geq 1-\alpha$. Note that in Theorem C.6, we assume non-empty lists, whereas in Proposition C.7 we allow empty prediction sets but apply the maximum operator $[x]^{+}:=\max\{x,0\}$. This is justified because the last term of Fano’s inequality relates to the probability of correct assignments $Y\in{\mathcal{C}}(X)$, which never happens for empty sets. See Proposition E.1 for the proof, where the same result appears.
@@
The bounds that we present in the main paper, model-based and simple Fano bounds, are actually derived through a slightly different root by leveraging the lower and upper bounds in the finite-sample guarantee of conformal prediction (Theorem 2.1). We derive these other two bounds in Appendix D.2.
C.5 Related Work on Information Theoretic Inequalities
@@
The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s—see for instance [2, 14, 42]. A key $f$-divergence inequality is the data processing inequality, which was used in information theory to establish various upper bounds on the achievability of coding schemes for different tasks—see [56, 55] for an extensive survey. Moreover, the data processing inequality for $f$-divergences provides a unified way of obtaining many classical and new results, for example Fano’s inequality [18]. The tightness of data processing inequalities is discussed in terms of Bregman’s divergence in [12, 32] and in terms of $\chi^{2}$-divergence in [55].
@@
When it comes to list decoding, there are a number of relevant inequalities in the literature. List decoding was introduced in context communication design by Elias [17] and Wozencraft [66]. For fixed list size, the information theoretic bounds on list decoding were obtained in [19] using error exponent analysis. A generalization of Fano’s inequality to list decoding was given in [15] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [51]. For fixed list size, stronger inequalities, some based on Arimoto-Rényi conditional entropy, were presented in [57]. Variable size list decoding was discussed in [55] using the notion of $E_{\gamma}$ resolvability first introduced in [50] related to the dependence testing bound. It was used again in the context of channel resolvability in [33], where some relevant inequalities have been obtained and discussed. A selection of the most relevant inequalities for list decoding can be found in [55].
Appendix D Proofs of Main Theoretical Results